# Diffusion-based Dynamic Contract for Federated AI Agent Construction in Mobile Metaverses

### v1

@@ Abstract
Mobile metaverses have attracted significant attention from both academia and industry, which are envisioned as the next-generation Internet, providing users with immersive and ubiquitous metaverse services through mobile devices. Driven by Large Language Models (LLMs) and Vision-Language Models (VLMs), Artificial Intelligence (AI) agents hold the potential to empower the creation, maintenance, and evolution of mobile metaverses. Currently, AI agents are primarily constructed using cloud-based LLMs and VLMs. However, several challenges hinder their effective implementation, including high service latency and potential sensitive data leakage during perception and processing. In this paper, we develop an edge-cloud collaboration-based federated AI agent construction framework in mobile metaverses. Specifically, Edge Servers (ESs), acting as agent infrastructures, collaboratively create agent modules in a distributed manner. The cloud server then integrates these modules into AI agents and deploys them at the edge, thereby enabling low-latency AI agent services for users. Considering that ESs may exhibit dynamic levels of willingness to participate in federated AI agent construction, we design a two-period dynamic contract model to continuously motivate ESs to participate in agent module creation, effectively addressing the dynamic information asymmetry between the cloud server and the ESs. Furthermore, we propose an Enhanced Diffusion Model-based Soft Actor-Critic (EDMSAC) algorithm to efficiently generate optimal dynamic contracts, in which dynamic structured pruning is applied to DM-based actor networks to enhance denoising efficiency and policy learning performance. Extensive simulations demonstrate the effectiveness and superiority of the EDMSAC algorithm and the proposed contract model.
Index Terms:
Mobile metaverses, AI agent construction, dynamic contract theory, enhanced diffusion models, deep reinforcement learning.
@@ I Introduction
With the rapid evolution of edge computing, metaverse, and Generative Artificial Intelligence (GenAI) technologies, mobile metaverses are emerging as a transformative digital ecosystem [1]. By integrating extended reality and advanced communication technologies, they establish interconnected virtual spaces accessible through mobile devices [2]. Positioned as a cornerstone of the next-generation Internet, mobile metaverses have attracted significant interest due to their robust levels of connectivity, enabling users to interact in real time with immersive virtual spaces. Furthermore, mobile metaverses hold the potential to revolutionize various fields, such as healthcare, education, and entertainment [2]. However, the development of mobile metaverses confronts several technical challenges. Firstly, conventional content creation methods often involve time-consuming and resource-intensive processes, leading to static virtual content that fails to dynamically adapt to real-time user interactions or environmental variations [3]. Secondly, the inherent limitations of mobile input modalities, such as touchscreen gestures and voice commands, would inevitably degrade the fluidity of immersive interactions.
@@
Currently, AI agent construction often relies on cloud-based LLMs and VLMs, which require the transmission of perceived data to remote servers for processing before returning the results [5]. Although AI agents hold significant potential in mobile metaverses, several critical challenges penalize their effective implementation: 1) The cloud-based AI agent architecture introduces inherent service latency, negatively impacting the user experience. For example, delays in processing and rendering real-time location data within augmented reality navigation systems may lead to user disorientation, operational inefficiencies, or even safety hazards in scenarios requiring precise spatial awareness [7]. 2) During peak usage periods, cloud servers may experience performance bottlenecks due to simultaneous service requests from large user bases [8]. Such bottlenecks result in response delays and potential service unavailability under high-load conditions, undermining the seamless interaction demanded by immersive metaverse environments. 3) The transmission of perceived data, which often contains sensitive personal information, remains vulnerable to interception during transmission and susceptible to breaches if stored centrally in cloud repositories [9]. This exposes users to risks such as unauthorized tracking, data misuse, and identity theft, particularly in mobile metaverses where data flows across heterogeneous networks.
@@
Motivated by the feasibility of federated model construction [10, 11, 12, 13, 14], we propose a novel edge-cloud collaboration-based federated AI agent construction framework, where Edge Servers (ESs) serve as agent infrastructures responsible for constructing agent modules, while the cloud server integrates these constructed modules into AI agents and eventually deploys them at the edge. Unlike traditional federated learning, which deals with the same task across participating nodes [15], the agent modules in the federated AI agent construction framework can be heterogeneous, with each module performing distinct functions. Nonetheless, in the federated AI agent construction framework, ESs exhibit different dynamic levels of Willingness to Participate (WTP) over multiple periods due to various factors, such as the high energy consumption of agent module construction and the prioritization of other computing tasks. Moreover, the WTP levels of ESs are typically unknown to the cloud server, leading to the intention of some ESs to free-ride on the efforts of others [16]. To this end, we design a dynamic contract-based incentive mechanism to incentivize ESs to participate sustainably in federated AI agent construction, addressing the challenge of dynamic information asymmetry. Particularly, we propose an enhanced diffusion-based Deep Reinforcement Learning (DRL) algorithm to identify optimal dynamic contracts. Our main contributions are summarized as follows:
@@
  • Federated AI Agent Construction Framework: We propose an edge-cloud collaboration-based federated AI agent construction framework in mobile metaverses. In the proposed framework, ESs serve as agent infrastructures, leveraging their computing, storage, and communication resources to create agent modules in a distributed manner. The cloud server then integrates these created agent modules into complete AI agents and deploys them at the edge, thereby providing low-latency AI agent services for users in mobile metaverses.
@@
  • Dynamic Contract-based Incentive Mechanism: Considering the dynamic WTP levels of ESs in federated AI agent construction, we design a two-period dynamic contract model to motivate ESs to continuously participate in agent module creation. The self-disclosure property of the proposed model ensures that each ES selects the contract best suited to its WTP level, thereby addressing the problem of dynamic information asymmetry. Notably, the two-period dynamic contract design maintains relatively low computational complexity and thus can be recursively extended to multiple periods.
@@
  • Diffusion Model for Dynamic Contract Design: We adopt an Enhanced Diffusion Model-based Soft Actor-Critic (EDMSAC) algorithm to identify optimal dynamic contracts. Specifically, the DM-based actor networks possess both analytical and generative capabilities, which can generate contract samples through an iterative denoising process. To enhance its denoising efficiency, we incorporate dynamic structured pruning techniques into the DM-based actor networks, enabling a more efficient and scalable exploration of the optimal dynamic contract policy. To the best of our knowledge, this is the first work to leverage DRL algorithms, especially diffusion-based DRL, for optimal dynamic contract design.
@@
  • Extensive Simulations for Performance Evaluation: We utilize LLMs called Qwen-VL [^1] https://github.com/QwenLM/Qwen-VL?tab=readme-ov-file to construct an AI agent capable of operating mobile phones to perform simple tasks. Specifically, we leverage Qwen-VL-Plus to create a perception module and Qwen-VL-Max to create decision-making, reflection, and planning modules. This construction process can be extended to the proposed framework. We then compare the EDMSAC algorithm with two DM-based DRL algorithms and three traditional DRL benchmark algorithms. Simulation results demonstrate the effectiveness of the proposed algorithm.
@@
The remainder of the paper is organized as follows: Section II reviews the related work. Section III introduces the edge-cloud collaboration-based federated AI agent construction framework. In Section IV, we propose the two-period dynamic contract model to motivate ESs to participate continuously in federated AI agent construction. Section V presents the EDMSAC algorithm for generating optimal dynamic contracts. In Section VI, we provide extensive simulation results to demonstrate the feasibility and effectiveness of the proposed framework and algorithm. Section VII concludes the paper.
@@ II-A Federated Model Construction
Federated Learning (FL) is a distributed Machine Learning (ML) framework that enables collaborative model training while preserving data privacy through secure aggregation, making it particularly suitable for applications with stringent privacy requirements [17]. Recent research has highlighted the effectiveness and potential of FL in constructing DT and ML models, particularly in enhancing privacy protection and optimizing resource utilization [10, 11, 12, 13, 14]. For instance, the authors in [11] combined FL with DT technology and introduced a trust mechanism to optimize the learning process. Similarly, the authors in [12] leveraged FL to optimize DT-enabled industrial Internet of Things systems. In [13], the authors proposed an FL architecture based on over-the-air computing [18], enabling the efficient and secure development of DT models. The authors in [14] proposed a generalized federated Reinforcement Learning (RL) framework that incorporates meta-learning techniques, which enables the fusion of RL models trained across multiple intelligent devices into a unified and generalizable model.
Combined with a device-edge-cloud collaborative computing architecture, FL can facilitate efficient model training by enabling small-scale learning on devices while leveraging the cloud for large-scale model updates. This decentralized approach minimizes network dependency, significantly enhancing the responsiveness and operational stability of AI agents and making it a promising solution for AI agent model construction. Inspired by the above works, we aim to develop a framework for federated AI agent construction.
II-B Contract Theory-based Applications
@@
As for the applications of multi-dimensional contracts [23, 24], the authors in [23] developed a multi-dimensional contract-theoretic framework to optimize the design of rewards of data in mobile networks under asymmetric information. In [24], the authors constructed a multi-dimensional contract theory-based framework, incorporating Prospect Theory (PT) and Generative DM (GDM)-based DRL algorithms to optimize the dynamic migration of embodied AI twins in vehicular embodied AI networks. As for the applications of dynamic contracts[22, 16], the authors in [16] proposed a two-period contract-based incentive mechanism to motivate users to participate in FL training under dynamic information asymmetry. In [22], the authors proposed a dynamic contract-based incentive mechanism to motivate vehicles to share their threat data for LLM fine-tuning. Inspired by the above works, we propose a dynamic contract-based incentive mechanism to incentivize ESs to participate in AI agent construction.
@@ II-C Deep Reinforcement Learning for Contract Design
DRL algorithms have been extensively used to identify optimization strategies, including optimal static contract design [25, 26, 27]. For instance, the authors in [25] employed Proximal Policy Optimization (PPO) algorithms to identify optimal contracts, incentivizing mobile devices to share data in mobile AI-Generated Content (AIGC) networks. In [26], the authors utilized the twin-delayed deep deterministic policy gradient algorithm to determine the incentive reward policy. Nevertheless, traditional DRL algorithms often struggle to directly identify optimal contracts due to the complex Individual Rationality (IR) and Incentive Compatibility (IC) constraints. Fortunately, GDMs excel at capturing high-dimensional information in dynamic and complicated network optimization scenarios [28]. Several studies have explored the use of GDMs for optimal static contract design [19, 24, 29]. For example, the authors in [19] proposed a DM-based DRL algorithm to generate optimal contracts under PT, addressing the problem of information asymmetry in edge AIGC service provision. However, existing works still rely on heuristic algorithms for optimal dynamic contract design[16, 22], which may not be practical in real scenarios. To fill this gap, we propose the EDMSAC algorithm to identify optimal dynamic contracts.
@@ III-A Framework Design
For federated AI agent construction, we consider a central cloud server and a set of geographically distributed ESs, denoted by $\mathcal{N}=\{1,\ldots,n,\ldots,N\}$, as shown in Fig. 1. During the whole AI agent construction period, AI agents can be split into multiple agent modules, such as perception, memory, and action, each responsible for specific tasks by processing the featured data perceived by ESs[5].
Before the beginning of agent module construction, the cloud server transmits the initial configuration parameters for the agent modules to the ESs (Step ①). The ESs, serving as agent infrastructures with sufficient resources (e.g., computing, storage, and bandwidth), leverage local LLMs and AI models (e.g., icon detection models) to construct the agent modules in a distributed manner (Step ②) [30]. Upon completing the construction, the ESs upload the constructed agent modules to the cloud server (Step ③). Finally, the cloud server creates the complete AI agents by integrating the uploaded agent modules (Step ④). Once constructed, the created AI agents can be deployed at the edge, allowing users to connect to them through the IP address of the current local area network with the ES [31].
Figure 1: The proposed edge-cloud collaboration-based federated AI agent construction framework. Note that the interface of personal assistants is from the OmAgent app [31] downloaded by our Android phone.
@@ III-B Model Quality of Agent Modules
Without loss of generality, we employ an $L$-Lipschitz continuous and $\varepsilon$-strongly convex loss function for the model training performed by ESs[32]. We denote $Q_{n}$ as the model quality of agent modules created by ES $n$. $Q_{n}$ is calculated from the number of training rounds $T_{n}$ that ES $n$ determines to take for creating agent modules, which is given by[32]
@@
where $\delta\in(0,\frac{2}{L})$, $L$, and $\varepsilon$ are pre-defined hyperparameters. Given that the second derivative of $Q_{n}$ in (1) is calculated as $\frac{\partial^{2}Q_{n}}{\partial(T_{n})^{2}}=-\big{[}\frac{(2-L\delta)\delta\varepsilon}{2}\ln 2\big{]}^{2}2^{\frac{-T_{n}(2-L\delta)\delta\varepsilon}{2}}<0$, $Q_{n}$ is a convex function with respect to $T_{n}$, indicating that there exists the optimal number of training rounds for ESs to obtain the highest model quality of agent modules.

### v2

@@ Abstract
Mobile metaverses are envisioned as a transformative digital ecosystem that delivers immersive, intelligent, and ubiquitous services through mobile devices. Driven by Large Language Models (LLMs) and Vision-Language Models (VLMs), Artificial Intelligence (AI) agents hold the potential to empower the creation, maintenance, and evolution of mobile metaverses, enabling seamless human-machine interaction and dynamic service adaptation. Currently, AI agents are primarily built upon cloud-based LLMs and VLMs. However, several challenges hinder their efficient deployment, including high service latency and a risk of sensitive data leakage during perception and processing. In this paper, we develop an edge-cloud collaboration-based federated AI agent construction framework in mobile metaverses. Specifically, Edge Servers (ESs), as agent infrastructures, first create agent modules in a distributed manner. The cloud server then integrates these modules into AI agents and deploys them at the edge, thereby enabling low-latency AI agent services for users. Considering that ESs may exhibit dynamic levels of willingness to participate in federated AI agent construction, we design a two-period dynamic contract model to continuously incentivize ESs to participate in agent module creation, effectively addressing the dynamic information asymmetry between the cloud server and ESs. Furthermore, we propose an Enhanced Diffusion Model-based Soft Actor-Critic (EDMSAC) algorithm to effectively generate optimal dynamic contracts. In the algorithm, we apply dynamic structured pruning to DM-based actor networks to enhance denoising efficiency and policy learning performance. Simulation results demonstrate that the EDMSAC algorithm outperforms the DMSAC algorithm by up to $23\%$ in optimal dynamic contract generation.
@@ I Introduction
With the advancement of extended reality, Generative Artificial Intelligence (GenAI), and wireless communication technologies, mobile metaverses are emerging as a transformative digital ecosystem [22]. Positioned as a cornerstone of the next-generation Internet, mobile metaverses have attracted significant attention from both academia and industry due to their robust levels of connectivity, enabling users to interact in real time with immersive and interconnected virtual spaces through mobile devices [34]. Furthermore, mobile metaverses hold the potential to revolutionize various fields, such as healthcare, education, and entertainment [34]. However, the development of mobile metaverses faces several technical challenges. Firstly, conventional content creation methods often involve time-consuming and resource-intensive processes, leading to static virtual content that fails to dynamically adapt to real-time user interactions or environmental variations [15]. Secondly, the inherent limitations of mobile input modalities, such as touchscreen gestures and voice commands, would inevitably degrade the fluidity of immersive interactions.
@@
Currently, AI agent construction often relies on cloud-based LLMs and VLMs, which require transmitting perceived data to remote servers for processing before returning results to users [6, 38]. Although AI agents hold significant promise for enabling mobile metaverses, their efficient implementation is hindered by several critical challenges: 1) The cloud-based AI agent architecture introduces inherent service latency, negatively impacting user experience. For example, latency in processing and rendering real-time location data within augmented reality navigation systems may lead to user disorientation, operational inefficiencies, or even safety hazards in scenarios requiring precise spatial awareness [24]. 2) During peak usage periods, cloud servers may experience performance bottlenecks due to simultaneous service requests from large user bases [19]. Such bottlenecks result in response latency and potential service unavailability under high-load conditions, undermining the seamless interaction demanded by immersive metaverse environments. 3) The transmission of perceived data, which often contains sensitive personal information, remains vulnerable to interception during transmission and susceptible to breaches if stored centrally in cloud repositories [18]. This exposes users to risks such as unauthorized tracking, data misuse, and identity theft, particularly in mobile metaverses where data flows across heterogeneous networks.
@@
Motivated by the feasibility of federated model construction [10, 25, 36, 9, 2], we propose a novel edge-cloud collaboration-based federated AI agent construction framework, where Edge Servers (ESs) with sufficient computing and storage resources serve as agent infrastructures responsible for constructing agent modules in a distributed manner, while the cloud server is responsible for integrating these constructed modules into AI agents and eventually deploys them at the edge. Unlike traditional federated learning, which deals with the same task across participating nodes [37], the agent modules in the federated AI agent construction framework can be heterogeneous, with each module performing distinct functions. Furthermore, whereas federated learning typically aims to jointly train a single global model until convergence [37], the construction of multiple agent modules can be done concurrently. Nonetheless, during the construction of agent modules, ESs exhibit different dynamic levels of Willingness to Participate (WTP) over multiple periods due to various factors [14], such as the high energy consumption of agent module construction and the prioritization of other computing tasks. Moreover, the WTP levels of ESs are typically unknown to the cloud server, leading to the intention of some ESs to free-ride on the efforts of others [14]. To this end, we design a dynamic contract-based incentive mechanism to incentivize ESs to participate sustainably in federated AI agent construction, addressing the challenge of dynamic information asymmetry.
Since dynamic contract design in practical scenarios typically involves changing network conditions [14], conventional optimization techniques that rely on complete prior knowledge of the environment must be frequently redesigned [29], which may not be applicable in practical scenarios. Fortunately, Deep Reinforcement Learning (DRL) algorithms have been widely utilized in dynamic scenarios, which can learn optimal strategies through repeated interactions with the environment [16]. Moreover, Generative Diffusion Models (GDMs) with a robust generative capability have shown a strong potential to improve DRL performance [5, 4, 28], particularly by mitigating the risk of convergence to suboptimal solutions. Therefore, we adopt an enhanced diffusion-based DRL algorithm to identify optimal dynamic contracts. The main contributions of this paper are summarized as follows:
@@
  • Federated AI Agent Construction Framework: We propose an edge-cloud collaboration-based federated AI agent construction framework in mobile metaverses. In the proposed framework, ESs serve as agent infrastructures, leveraging their computing, storage, and communication resources to construct agent modules in a distributed manner. The cloud server then integrates these created agent modules into complete AI agents and deploys them at the edge, thereby providing low-latency AI agent services for users in mobile metaverses.
@@
  • Dynamic Contract-based Incentive Mechanism: Considering the dynamic WTP levels of ESs in federated AI agent construction, we propose a two-period dynamic contract model to motivate ESs to continuously participate in agent module creation. The self-disclosure property of the proposed model ensures that each ES selects the contract best suited to its WTP level, thereby addressing the problem of dynamic information asymmetry. Notably, the two-period dynamic contract design maintains relatively low computational complexity and thus can be recursively extended to multiple periods.
@@
  • Diffusion Models for Dynamic Contract Design: We adopt an Enhanced DM-based Soft Actor-Critic (EDMSAC) algorithm to identify optimal dynamic contracts. Specifically, DM-based actor networks possess both analytical and generative capabilities, which can generate dynamic contract samples through an iterative denoising process. To enhance denoising efficiency, we incorporate dynamic structured pruning techniques into DM-based actor networks, enabling a more efficient and scalable exploration of the optimal dynamic contract policy. To the best of our knowledge, this is the first work to leverage DRL algorithms, especially diffusion-based DRL, for optimal dynamic contract design.
@@
  • Extensive Simulations for Performance Evaluation: We utilize LLMs called Qwen-VL [^1] https://github.com/QwenLM/Qwen-VL?tab=readme-ov-file to construct an AI agent capable of operating mobile phones to perform simple tasks. Specifically, we leverage Qwen-VL-Plus to create a perception module and Qwen-VL-Max to create decision-making, reflection, and planning modules. This construction process can be extended to the proposed framework. We then compare the EDMSAC algorithm with two DM-based DRL algorithms and three representative DRL benchmark algorithms. Simulation results demonstrate that the EDMSAC algorithm can achieve up to a $23\%$ improvement over the DMSAC algorithm [4] in optimal dynamic contract generation.
@@
The remainder of the paper is organized as follows: Section II reviews the related work, where Table I summarizes the comparison of the current literature and this paper in optimal contract design. Section III introduces the edge-cloud collaboration-based federated AI agent construction framework. In Section IV, we propose the two-period dynamic contract model to motivate ESs to continuously participate in federated AI agent construction. Section V presents the EDMSAC algorithm for generating optimal dynamic contracts. In Section VI, we provide extensive simulation results to demonstrate the feasibility and effectiveness of the proposed framework and algorithm. Section VII concludes the paper.
TABLE I: Comparison Between the Current Literature and This Paper in Optimal Contract Design
Literature [10] [14] [29] [16] [31] [27] [12] [7] [33] [41] [32] [40] Our Paper
Properties Individual Rationality
Incentive Compatibility
Intertemporal Consistency
Solutions Deep Reinforcement Learning
Diffusion Models
Enhanced Diffusion Models
@@ II-A Federated Model Construction
Federated Learning (FL) is a distributed Machine Learning (ML) framework that enables collaborative model training while preserving data privacy through secure aggregation, making it particularly suitable for applications with stringent privacy requirements [23]. Recent research has highlighted the effectiveness and potential of FL in constructing DT and ML models, particularly in enhancing privacy protection and optimizing resource utilization [10, 25, 36, 9, 2]. For instance, the authors in [25] combined FL with DT technology and introduced a trust mechanism to optimize the learning process. Similarly, the authors in [36] leveraged FL to optimize DT-enabled industrial Internet of Things systems. In [9], the authors proposed an FL architecture based on over-the-air computing, enabling the efficient and secure development of DT models. In [2], the authors proposed a generalized federated Reinforcement Learning (RL) framework that incorporates meta-learning techniques, which enables the fusion of RL models trained across multiple intelligent devices into a unified and generalizable model.
Combined with a device-edge-cloud collaborative computing architecture, FL can facilitate efficient model training by enabling small-scale learning on devices while leveraging the cloud for large-scale model updates. This decentralized approach minimizes network dependency, significantly enhancing the responsiveness and operational stability of AI agents and making it a promising solution for AI agent model construction. To the best of our knowledge, there is currently no existing study on the construction of AI agents in wireless mobile environments. Motivated by the aforementioned studies, we aim to develop a novel federated AI agent construction framework, exploring the feasibility and potential of mobile AI agents in wireless networks.
II-B Contract Theory-based Wireless Applications
@@
As for the applications of multi-dimensional contracts [33, 41], the authors in [33] developed a multi-dimensional contract model to optimize the data rewarding design in mobile networks under asymmetric information. In [41], the authors constructed a multi-dimensional contract theory-based framework, incorporating Prospect Theory (PT) and GDM-based DRL algorithms to optimize the dynamic migration of embodied AI twins in vehicular embodied AI networks. As for the applications of dynamic contracts [7, 14], the authors in [14] proposed a two-period contract-based incentive mechanism to motivate users to participate in FL training under dynamic information asymmetry. In [7], the authors proposed a dynamic contract-based incentive mechanism to motivate vehicles to share their threat data for LLM fine-tuning. Inspired by the above works, we propose a dynamic contract-based incentive mechanism to motivate ESs to continuously participate in AI agent construction.
@@ II-C Deep Reinforcement Learning for Contract Design
RL algorithms have been extensively used to identify optimization strategies [21, 20], including optimal static contract design [32, 40, 16]. For instance, the authors in [32] employed Proximal Policy Optimization (PPO) algorithms to identify optimal contracts, incentivizing mobile devices to share data in mobile AI-Generated Content (AIGC) networks. In [40], the authors utilized the twin-delayed Deep Deterministic Policy Gradient (DDPG) algorithm to determine the incentive reward policy. Nevertheless, traditional DRL algorithms often struggle to directly identify optimal contracts due to the complex Individual Rationality (IR) and Incentive Compatibility (IC) constraints. Fortunately, GDMs excel at capturing high-dimensional information in dynamic and complicated network optimization scenarios [5]. Several studies have explored the use of GDMs for optimal static contract design [31, 41, 29]. For example, the authors in [31] proposed a GDM-based DRL algorithm to generate optimal contracts under PT, addressing the problem of information asymmetry in edge AIGC service provision. However, existing studies still rely on heuristic algorithms for optimal dynamic contract design [14, 7], which may not be practical in real scenarios. To fill this gap, we propose the EDMSAC algorithm to identify optimal dynamic contracts.
@@ III-A Framework Design
For federated AI agent construction, we consider a central cloud server and a set of geographically distributed ESs, which are usually installed in cellular base stations or roadside units [41], denoted by $\mathcal{N}=\{1,\ldots,n,\ldots,N\}$. These ESs possess computing, storage, and communication capabilities sufficient to construct, cache, and transmit agent modules [1]. During the AI agent construction period, AI agents can be split into multiple agent modules, such as perception, memory, and action [6, 38], each responsible for specific tasks based on the featured data perceived by ESs.
As shown in Fig. 1, before the beginning of agent module construction, the cloud server transmits the initial configuration requirements of agent modules to the ESs (Step ①) [10]. The ESs, serving as agent infrastructures with sufficient resources such as computing, storage, and communication, leverage local LLMs and AI models (e.g., icon detection models) to construct the agent modules in a distributed manner (Step ②) [26]. Upon completing the construction, the ESs upload the constructed agent modules to the cloud server (Step ③) [10]. Finally, the cloud server creates the complete AI agents by integrating the uploaded agent modules, i.e., connecting and harmonizing these agent modules into a complete agent (Step ④) [1]. Once constructed, the AI agents can be deployed at the edge, allowing users to connect to them through the Internet Protocol (IP) address of the current local area network with the ESs [39], facilitating low-latency AI agent services.
Figure 1: The proposed edge-cloud collaboration-based federated AI agent construction framework. Note that the interface of personal assistants is from the OmAgent application [39] downloaded by our Android phone.
@@ III-B Model Quality of Agent Modules
Agent module construction involves fine-tuning pre-trained large models and other AI models to better process downstream tasks [6]. For example, during the perception module construction, ESs are required to fine-tune visual perception models, such as vision transformer models, to extract corresponding semantic representations from visual inputs using samples of feature data. The extracted semantic content is then passed to LLMs for subsequent processing and reasoning. To evaluate the performance of agent module construction, we employ an $L$-Lipschitz continuous and $\varepsilon$-strongly convex loss function to approximately calculate the model quality of agent modules [1, 13], which is a widely used definition of model accuracy in federated learning [35, 1]. We denote $Q_{n}$ as the model quality of agent modules created by ES $n$, which is calculated from the number of training rounds $T_{n}$ that ES $n$ determines to take for creating agent modules, as given by [35]
@@
where $\delta\in(0,\frac{2}{L})$, $L$, and $\varepsilon$ are pre-defined hyperparameters [1]. Given that the second derivative of $Q_{n}$ in (1) is calculated as $\frac{\partial^{2}Q_{n}}{\partial(T_{n})^{2}}=-\big[\frac{(2-L\delta)\delta\varepsilon}{2}\ln 2\big]^{2}2^{\frac{-T_{n}(2-L\delta)\delta\varepsilon}{2}}<0$, $Q_{n}$ is a convex function with respect to $T_{n}$, indicating that there exists the optimal number of training rounds for ESs to obtain the highest model quality of agent modules.